### v1

# DRPO: Efficient Reasoning via Decoupled Reward Policy Optimization

Abstract
Recent large reasoning models (LRMs) driven by reinforcement learning algorithms (e.g., GRPO) have achieved remarkable performance on challenging reasoning tasks. However, these models suffer from overthinking, generating unnecessarily long and redundant reasoning even for simple questions, which substantially increases computational cost and response latency. While existing methods incorporate length rewards to GRPO to promote concise reasoning, they incur significant performance degradation. We identify the root cause: when rewards for correct but long rollouts are penalized, GRPO’s group-relative advantage function can assign them negative advantages, actively discouraging valid reasoning. To overcome this, we propose Decoupled Reward Policy Optimization (DRPO), a novel framework that decouples the length-based learning signal of correct rollouts from incorrect ones. DRPO ensures that reward signals for correct rollouts are normalized solely within the positive group, shielding them from interference by negative samples. The DRPO’s objective is grounded in integrating an optimized positive data distribution, which maximizes length-based rewards under a KL regularization, into a discriminative objective. We derive a closed-form solution for this distribution, enabling efficient computation of the objective and its gradients using only on-policy data and importance weighting. Of independent interest, this formulation is general and can incorporate other preference rewards of positive data beyond length. Experiments on mathematical reasoning tasks demonstrate DRPO’s significant superiority over six efficient reasoning baselines. Notably, with a 1.5B model, our method achieves 77% length reduction with only 1.1% performance loss on simple questions like GSM8k dataset, while the follow-up baseline sacrifices 4.3% for 68% length reduction. The code is available at https://github.com/Optimization-AI/DRPO
1 Introduction
Recently, large reasoning models (LRMs) driven by Reinforcement Learning (RL) (Guo et al., 2025; Team et al., 2025b) have demonstrated remarkable performance on complex reasoning tasks like mathematics, coding, and scientific problem-solving. Unlike conventional language models that focus on direct thoughts and solutions, LRMs improve performance by generating extended chain-of-thought paths (Wei et al., 2022), allowing them to revisit intermediate steps, correct errors, and even explore alternative reasoning paths. This approach equips LRMs with stronger reasoning abilities and has become a standard paradigm to develop models capable of solving complex tasks.
However, existing LRMs suffer from overthinking with lengthy and redundant reasoning paths. As demonstrated by Chen et al. (2024), reasoning models like DeepSeek-R1 (Guo et al., 2025) need to generate about 1,000 tokens to answer “what is the answer of 2 plus 3”, while only around 10 tokens are needed for non-reasoning models. Such overly generated reasoning paths raise significant issues, leading to substantially increased computational cost and longer inference time. Numerous studies have been conducted to explore ways to eliminate redundant reasoning and improve the reasoning efficiency. A popular strategy is to introduce explicit reward shaping with length penalties in RL to guide the model toward concise reasoning (Arora and Zanette, ; Huang et al., ; Xiang et al., ; Aggarwal and Welleck, ), e.g., penalizing rewards of correct answers based on reasoning length to encourage shorter reasoning. Nevertheless, almost all existing methods fall short in preserving performance while shortening reasoning, causing substantial performance loss. This raises a key question: how to guide RL for efficient reasoning with minimal performance drop?
We identify the root cause underlying the insufficiency of existing RL-based efficient reasoning methods. The recent advancement of training efficient reasoning models has been largely built upon the Group Relative Policy Optimization (GRPO) framework, due to its groundbreaking performance (Guo et al., 2025). GRPO’s effectiveness hinges on its group-relative advantage function, which normalizes a rollout’s reward against the group average to create a learning signal that distinguishes positive from negative examples. Yet, this very strength becomes its greatest weakness when moving beyond simple correctness. We demonstrate that the framework is fundamentally ill-suited for composite rewards. Incorporating a length penalty reduces the reward for correct but long answers, often pushing their group-relative advantage below zero. Consequently, GRPO is misled into interpreting verbose correct answers as negative examples, discouraging valid reasoning and creating a significant optimization barrier (Figure 1).
An effective mechanism must not only distinguish right from wrong but also efficient right from inefficient right—assigning a strong positive signal to concise answers and a weaker positive signal to verbose ones, all while suppressing incorrect reasoning. To this end, we introduce Decoupled Reward Policy Optimization (DRPO), a novel RL framework that fundamentally rethinks how learning signals are constructed. DRPO’s core innovation is the decoupling of the learning signal calculation: it normalizes rewards for a correct rollout only against other correct rollouts, completely insulating them from the negative examples that corrupt GRPO’s signal. This ensures that length penalty proportionally reduces the positive signal of a long correct answer but never pushes it into negative territory, thereby achieving a more favorable trade-off between efficiency and accuracy. We formalize this intuition by deriving a generalized objective in a discriminative RL framework. This objective integrates a perturbed version of the on-policy positive data distribution, where the perturbation is designed explicitly to maximize a length-based reward. We derive a closed-form solution of the perturbed distribution, which allows us to efficiently optimize the objective without any additional data collection, using only on-policy samples via importance weighting.
Our contributions are four-fold:
  • We diagnose a critical, previously overlooked deficiency in the widely-adopted GRPO framework: its group-relative advantage function is ill-suited for correctness–length composite rewards and actively harms learning when incentivizing efficiency.
  • We propose Decoupled Reward Policy Optimization (DRPO), a new paradigm that decouples learning signals for positive and negative data. DRPO provides consistent, uncorrupted policy gradients for multi-reward optimization (e.g., correctness and length).
  • We derive a rigorous formulation for DRPO by integrating a reward-maximizing, perturbed positive data distribution directly into a discriminative objective. We obtain a tractable closed-form solution to the perturbed distribution, yielding a practical algorithm requiring only on-policy data with no overhead.
  • We conduct experiments to demonstrate the superiority of DRPO in training efficient reasoning models, substantially outperforming strong baselines across different model sizes and various mathematical reasoning benchmarks.
2 Related work
Large Reasoning Models. Earlier structured prompting approaches such as Chain-of-Thought (CoT) (Wei et al., 2022), Tree-of-Thought (ToT) (Yao et al., 2023), and Graph-of-Thought (GoT) (Besta et al., 2024) demonstrated the importance of decomposing complex problems into intermediate steps. However, these methods rely heavily on prompting and search heuristics, lacking a unified learning framework to optimize reasoning efficiency and robustness.
The breakthrough came with DeepSeek-R1, which revealed that large reasoning models (LRMs) trained via large-scale reinforcement learning, particularly GRPO (Shao et al., 2024), can autonomously acquire advanced reasoning behaviors such as branching, verification, and backtracking. This success inspired numerous follow-up studies: some aimed at reproducing the effectiveness of GRPO (Wen et al., 2025; Luo et al., 2025b; He et al., 2025), while others investigated its limitations and proposed refinements to further enhance reasoning performance (Liu et al., 2025c; Yu et al., 2025; Li et al., 2025a). In parallel, many open-weight reasoning models have adopted GRPO, including Qwen-3 (Yang et al., 2025a), GLM-4.5 (Team et al., 2025a), K2-Think (Cheng et al., 2025), and Goedel-solver (Lin et al., 2025), among others. Despite these advances, existing LRMs often suffer from overthinking—producing unnecessarily long and redundant reasoning even for simple problems. This work aims to directly address this issue by training LRMs to reason both efficiently and effectively.
Efficient Reasoning in LRMs. To address the issue of overthinking in LRMs, a variety of techniques have been proposed (Sui et al., 2025; Yue et al., 2025), including (1) training-free methods, which shorten the reasoning paths via prompt (Aytes et al., 2025; Han et al., 2024) or manipulating the decoding process (Yang et al., 2025b; Yong et al., 2025; Wang et al., 2025a; Liu et al., 2025a; Wang et al., 2025b); (2) Supervised Fine-tuning (SFT) methods, which rely on compressed reasoning datasets for finetuning. These datasets are curated via token-level selection (Yuan et al., 2025; Xia et al., 2025; Zhuang et al., 2025), step-level selection (Xiao et al., 2025; Cui et al., 2025; Wang et al., 2025c), path-level selection (Munkhbat et al., 2025; Ghosal et al., 2025); (3) RL-based methods, which carefully design reward mechanisms to guide the model to reason efficiently (Hou et al., ; Aggarwal and Welleck, ; Arora and Zanette, ; Xiang et al., ; Luo et al., ; Huang et al., ; Yi et al., 2025; Liu et al., 2025b; Xu et al., 2025; Fang et al., 2025; Li et al., 2025b).
Among these techniques, RL-based methods have been demonstrated to be one of the most effective approaches due to their scalability and flexibility. Specifically, L1 (Aggarwal and Welleck, ) integrates a length constraint into the reward function to encourage higher performance while meeting the length goal specified in the prompt. Arora and Zanette employs online RL with a length penalty based on the distribution of correct answers, penalizing correct responses longer than the average while encouraging those shorter than the average. ALP (Xiang et al., ) adaptively adjusts length penalties according to problem difficulty, measured by pass rate, assigning stronger penalties to high pass-rate problems to discourage overlong reasoning. HAPO (Huang et al., ) keeps track of the minimum length of correct responses for each question, penalizing outputs that exceed this length while rewarding those that are shorter. Nevertheless, these methods all suffer from misleading learning signals and fall short in preserving performance while shortening reasoning, due to the limitation of their adopted relative advantage function.
Figure 1: Illustration of the limitation of GRPO with length penalty and the benefit of our approach. Suppose $[1,1,1,0,0,0]$ are the accuracy rewards for 6 responses, and $[0.73,0.6,0.2,0,0,0]$ are the rewards after applying the length penalty to correct answers. Using the group-relative advantage calculation of GRPO, the advantages for the third response shift from 1 (without length penalty) to -0.17 (with length penalty added), inadvertently penalizing the third correct response, which may substantially harm performance. In contrast, our proposed DRPO reduces the learning signal for lengthy and correct responses but never pushes them to the negative territory.
3 Limitation of incorporating length penalty into GRPO
Notations. We study the fine-tuning of a generative reasoning model $\pi_{\theta}$ parameterized by $\theta$. At each learning step, the previous model is denoted by $\pi_{\text{old}}$, which is responsible for generating answers to a given set of questions. For a question $q\in\Sigma^{*}$ (including its prompt), the output $o\in\Sigma^{*}$ is sampled from $\pi_{\text{old}}(\cdot|q)$, consisting of both reasoning traces and the final answer, where $\Sigma^{*}$ denotes the space of all sequences of tokens with arbitrary length. More concretely, $o$ is generated sequentially at the token level: $o_{t}\sim\pi_{\text{old}}(\cdot|q,o_{<t})$, for $t=1,\cdots,|o|$. The correctness reward $r_{c}(o|q)\in\{1,0\}$ for a given question $q$ and its corresponding answer in the output $o$ is verified by either matching the extracted answer against the ground-true answer or a formal verification tool (Guo et al., 2025; Zhang et al., 2025). Let $\pi_{\text{old}}^{+}(\cdot|q)$ denote the conditional distribution of outputs when the reward is one (i.e., correct answers) and $\pi_{\text{old}}^{-}(\cdot|q)$ denote the conditional distribution of outputs when the reward is zero (i.e., incorrect answers). Let $\cdot]_{+}=\max(\cdot,0)$ denote the hinge function.
Following the success of DeepSeek-R1, Group Relative Policy Optimization (GRPO) is widely adopted in existing RL-based efficient reasoning methods to estimate the relative advantage from group rewards instead of the critic model. In the following, we illustrate the limitations of incorporating length penalty into GRPO for promoting efficient reasoning. We note that similar limitations occur to other RL methods, such as RLOO (Ahmadian et al., 2024), and other REINFORCE-based methods (Hu et al., 2025; Chu et al., 2025), which couple rewards for correct and incorrect answers to compute advantages for learning.
The GRPO objective for maximization is given by:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q}\mathbb{E}_{\{o_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|q)}$ (1)
$\bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big(r_{i,t}A(o_{i}|q),\text{clip}(r_{i,t},1-\epsilon,1+\epsilon)A(o_{i}|q)\Big)\bigg]-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$
where $r_{i,t}=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\text{old}}(o_{i,t}|q,o_{i,<t})}$, $\pi_{\text{ref}}$ is a frozen reference model, $\mathbb{D}_{\text{KL}}(\cdot,\cdot)$ denotes the KL divergence between two distributions, and $A(o_{i}|q)$ denotes the relative advantage of output $o_{i}$, which quantifies how much better the reward of $o_{i}$ denoted by $r(o_{i}|q)$ compared to the group average reward and guides the learning direction. Specifically, $A(o_{i}|q)$ is computed by
$A(o_{i}|q)=\frac{r(o_{i}|q)-\text{mean}({r(o_{1}|q),r(o_{2}|q),\cdots,r(o_{G}|q)})}{\text{std}({r(o_{1}|q),r(o_{2}|q),\cdots,r(o_{G}|q)})}.$ (2)
Let us consider incorporating a length penalty into GRPO. Existing RL methods with length control reveal a shared principle: penalizing rewards of correct answers based on reasoning length to encourage shorter reasoning, e.g., $r(o|q)=r_{c}(o|q)-r_{l}(o|q)$, where $r_{l}(\cdot)$ is a length-based cost or reward function (refer to Table 2 for detailed formulations). The consequence is that the reward of a correct with long output will be shifted down relatively. Incorporating this reward into GRPO’s group relative advantage calculation may bias the intended effect, misleading the learning process. Let us consider an illustrative example in Figure 1. Suppose there are six generated outputs with different lengths, whose correctness rewards are $[1,1,1,0,0,0]$ and corresponding lengths are $[2000,2500,4000,2800,3800,3200]$. After combining the length reward with correctness reward, the combined reward of each answer becomes $[0.73,0.6,0.2,0,0,0]$  [^1] These values are calculated with the formula proposed in  (Arora and Zanette, ). We note that various reward combination designs in the literature lead to the same issue. Refer to Appendix A.3 for a detailed discussion. . So far, it looks like that all the designs work well since short correct answers have larger positive rewards and longer correct answers have smaller positive rewards while incorrect answers have zero rewards. However, when computing group relative advantage $A(o|q)$ in GRPO (i.e., Eqn. (2)), the advantage for the third correct response shifts from 1 (without length penalty) to -0.17 (with length penalty added). This negative signal will discourage valid reasoning and create a significant optimization barrier, which may substantially harm performance.
4 DRPO: Decoupled Reward Policy Optimization
The main reason of getting a negative learning signal in GRPO for a verbose correct answer is that its reshaped reward could become less than the mean reward of all samples including positive and negative ones. Our approach to avoid this issue is to decouple the rewarding of positive and negative samples so that the length rewards are only normalized within the positive group. To this end, we develop our approach based on a recent work (Li et al., 2025a), which proposes a discriminative optimization framework (DisCO) that directly increases the generative likelihood of positive answers and decrease that of negative answers. Below, we first introduce DisCO and then present our novel approach of integrating length rewards into DisCO’s objective.
4.1 Discriminative Constrained Policy Optimization (DisCO)
DisCO was proposed to address several inherent limitations of GRPO, including difficulty bias and clipping operations. Let $s_{\theta}(o,q)$ be a scoring function, which measures the generative likelihood of answer $o$ given the input $q$. In this paper, we will consider $s_{\theta}(o,q)=\frac{1}{|o|}\sum_{t=1}^{|o|}\log\pi_{\theta}(o_{t}|q,o_{<t})$, which is effective as demonstrated in (Li et al., 2025a). The objective of DisCO is formulated as:
$\max_{\theta}\quad\mathbb{E}_{q}\left[\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}s_{\theta}(o,q)-\tau\log\bigg(\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\exp\bigg(\frac{s_{\theta}(o^{\prime},q)}{\tau}\bigg)\bigg)\right],$ (3)
$s.t.\quad\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})\leq\delta,$
where $\delta>0$ is a hyper-parameter. The intuition behind this formulation is straightforward: it increases the scores of positive responses $o\sim\pi_{\text{old}}^{+}(\cdot|q)$ while decreasing the scores of negative responses, aggregated through a log-sum-exp function. The log-sum-exp has its roots in discriminative learning, appearing in losses such as cross-entropy and contrastive loss, and naturally emphasizes hard negatives by assigning them larger learning signal. The constraint $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})\leq\delta$, inspired by TRPO (Schulman et al., 2015), is added to ensure the stability of training.
While DisCO demonstrates impressive gains in reasoning performance over GRPO, the length of its reasoning is uncontrolled, leaving the challenge of enhancing reasoning efficiency unresolved. Moreover, the above objective is derived under a binary reward setting, which does not accept flexible reward design. In the following, we discuss how to incorporate length rewards into the framework to encourage efficient reasoning.
4.2 Decoupled Reward Policy Optimization
We consider a simple length reward $r_{l}(o)=1-\frac{|o|}{C}$ for any correct response $o$, where $C$ is a constant denoting maximum response length. An intuitive idea is to assign a weight to positive answers before their scores $s_{\theta}(o,q)$ in (3) such that a shorter answer is assigned with a larger weight than a longer answer. Below, we formalize this idea by proposing a principled objective.
Our goal is to maximize the score of correct outputs with high length rewards while penalizing those of wrong outputs regardless of their lengths. Suppose we have a distribution $P_{q}^{*}$, which specifies a distribution of correct outputs with high length rewards given a question $q$. Then, we can modify the objective in (3) as
$\max\mathbb{E}_{q}\left[\mathbb{E}_{o\sim P_{q}^{*}}s_{\theta}(o,q)-\tau\log\bigg(\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\exp\bigg(\frac{s_{\theta}(o^{\prime},q)}{\tau}\bigg)\bigg)\right].$ (4)
This can be explained that if we have an off-policy data distribution $P^{*}_{q}$ of correct outputs with high length rewards, we can use its sampled data to steer the model training such that it generates correct outputs with high length rewards more likely. However, the issue is that $P^{*}_{q}$ is not readily available. Although a naive solution is to curate such data manually, as in SFT-based efficient reasoning methods, it requires substantial human effort and lacks scalability. To address this, we draw insights from reinforcement learning from human feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022; Ziegler et al., 2019) by finding an optimal policy $P^{*}_{q}$ that maximizes the length reward with a KL regularization:
$P_{q}^{*}=\arg\max_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}r_{l}(o)-\lambda\mathbb{D}_{\text{KL}}(P,\pi_{\text{old}}^{+}(\cdot|q)),$ (5)
where $\lambda>0$ is a regularization parameter, $\mathcal{P}$ denotes the set of all probability measures $P$ on correct data given $q$, which are absolutely continuous with respect to $\pi_{\text{old}}^{+}(\cdot|q)$, i.e., $\pi_{\text{old}}^{+}(o|q)=0$ indicates $P(o)=0$.
However, unlike RLHF that uses a LLM to learn $P^{*}$, we derive its closed analytical solution similar to (Rafailov et al., 2023) (See Appendix A.1 for a complete derivation):
$P_{q}^{*}(o)=\frac{\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}.$ (6)
As a result, we have
$\mathbb{E}_{o\sim P_{q}^{*}}s_{\theta}(o,q)$ $=\sum_{o\in\Sigma^{*}}\frac{\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}s_{\theta}(o,q)$
$=\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(o|q)}\frac{\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}s_{\theta}(o,q).$
Plugging this formulation back into (4), we obtain the final objective function:
$\max\quad\ \mathbb{E}_{q}\left[\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\frac{\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}s_{\theta}(o,q)-\tau\log\bigg(\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\exp\bigg(\frac{s_{\theta}(o^{\prime},q)}{\tau}\bigg)\bigg)\right]$ (7)
$s.t.\quad\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})\leq\delta.$
It is notable that the final objective only relies on the on-policy data, and it has an explanation that each positive data is assigned with a weight $\omega(o|q)=\frac{\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}$ informed by its length but normalized only withing the positive data. It is notable that when $\lambda=+\infty$, then $\omega(o|q)=1$ and the above objective reduces to that of DisCO in (3).
We solve the optimization problem (7) similarly as (Li et al., 2025a). In particular, the expectations are replaced by empirical averages and the KL divergence is estimated by using sampled data, and the constraint is handled by adding a penalty function $\beta_{0}[\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})-\delta]_{+}^{2}$ to the objective, where $\beta_{0}$ is a penalty constant. For completeness, we present a full algorithm for solving (7) in Algorithm 1 in the Appendix. We refer to this method as Decoupled Reward Policy Optimization (DRPO).
5 Experiments
Datasets. We validate our method on mathematical reasoning tasks. Specifically, we train models on the DeepScaleR-Preview-Dataset (Luo et al., 2025c), which consists of approximately 40.3k question-answer pairs sourced from AIME problems from 1984 to 2023, AMC problems before 2023, Omni-MATH (Gao et al., 2024) and Still (Min et al., 2024) datasets. We evaluate all the models on math problems with different levels of difficulty, including (a) easy level: GSM8K (Cobbe et al., 2021), (b) medium level: MATH-500 (Hendrycks et al., 2021), (c) hard level: OlympiadBench (He et al., 2024), and (d) very hard level: AIME (aggregating 2024 and 2025).
Models. We adopt two reasoning models as our base models: DeepSeek-R1-Distill-Qwen-1.5B model and DeepSeek-R1-Distill-Qwen-7B, and conduct RL fine-tuning from them.
Baselines. We compare our methods with six of the most recent state-of-the-art efficient reasoning methods, including (1) the method in (Arora and Zanette, ), which integrates length reshaped rewards into the RLOO advantage function, referred to as RLOO-LP; (2) ALP (Xiang et al., ), which uses a length penalty in GRPO that is the length scaled by the solving rate of each question; (3) HAPO (Huang et al., ), which penalizes the responses longer than the shortest correct answer in the history while rewarding those that are shorter; (4) L1-max (Aggarwal and Welleck, ), which is a reasoning language model that produces outputs satisfying a maximum length constraint given in the prompt; (5) ShorterBetter (SB) (Yi et al., 2025), which aims to match Sample Optimal Length defined as the shortest correct response among multiple generations; (6) LASER-D (Liu et al., 2025b), which employs a step length reward function based on difficulty-aware dynamic target length. We train models for methods (1)-(3) using the experimental settings described below. For methods (4)-(6), we evaluate the models provided in their original works. All the compared models were finetuned from the same base models on the DeepScaleR-Preview dataset, except L1-max, which was trained on DeepScaleR-1.5B-Preview. We summarize different reward designs of the above baselines in Table 2 in Appendix.
Training Details. For all the training, we employ the AdamW optimizer with a weight decay of 0.01 and set the learning rate to a constant $2\mathrm{e}^{-6}$ for 1.5B model and $1\mathrm{e}^{-6}$ for 7B model, following  Li et al. (2025a). We set the batch size to 128 for each step of RL, the mini-batch size to 32 for each iteration of model update, and sample 8 responses per question for training. For RLOO-LP, we tune their weight parameter $\alpha\in\{0.05,0.1,0.2\}$. For ALP, we tune their penalty weight $\beta\in\{1\mathrm{e}^{-9},1\mathrm{e}^{-8},1\mathrm{e}^{-7}\}$. For HAPO, we tune their weight parameter $w\in\{0.01,0.1,1\}$. For the proposed method, we tune $\lambda\in\{0.5,0.2,0.1\}$. These parameters serve the same role that controls the tradeoff between efficiency and accuracy. For all other hyperparameters, we follow the default values from their official papers. Details are provided in Appendix A.2. The generation budget is limited to 8k tokens for both training and evaluation.
Evaluation. We use Pass@1 averaged over the 16 generated answers per prompt as the performance metric and use the averaged number of tokens as the length metric. For all methods, we train the model for 1000 RL steps to enable convergence and conduct evaluation every 200 steps. The models with the best pass@1 are reported, as we aim to enhance reasoning efficiency with minimal performance reduction. For models that are trained by us, we set temperature = 0.6 and top-p = 0.95, consistent with the training setup. For L1-MAX and LASER-D, we also use temperature = 0.6 and top-p = 0.95, while we use temperature = 0.9 and top-p = 0.9 for SB, following the original paper.
In addition to pass@1 and reasoning length, following (Luo et al., 2025a; Yi et al., 2025), we also adopt Accuracy Efficiency Score (AES) as a supplementary metric. AES integrates performance and reasoning length into a single measure, directly quantifying the trade-off between accuracy and computational cost. The AES is computed by:
$\text{AES}=\begin{cases}\alpha*\Delta_{\text{Length}}+\beta*|\Delta_{\text{Acc}}|,&\text{if }\Delta_{\text{Acc}}\geq 0,\\ \alpha*\Delta_{\text{Length}}-\gamma*|\Delta_{\text{Acc}}|,&\text{if }\Delta_{\text{Acc}}<0.\end{cases}$
where $\alpha,\beta,\gamma>0$, $\Delta_{\text{Length}}=\frac{\text{Length}_{\text{ref}}-\text{Length}_{\text{model}}}{\text{Length}_{\text{ref}}}$ and $\Delta_{\text{Acc}}=\frac{\text{Acc}_{\text{model}}-\text{Acc}_{\text{ref}}}{\text{Acc}_{\text{ref}}}$, where the quantities with subscript “ref" means each method’s baseline that does not consider length reward, and that with subscript “model" means the model for evaluation. In our experiments, we use the default values $\alpha=1,\beta=3$ same as (Luo et al., 2025a), but set $\gamma=10$ to emphasize the importance of minimizing performance degradation.
Figure 2: Training dynamics of DRPO with different regularization weights $\lambda$. The left two plots are for fine-tuning the 1.5B model, and the right two are for fine-tuning the 7B model. $\lambda=+\infty$ denotes the reference method DisCO, which does not incorporate length rewards in training.
5.1 Visualization of Learning Process
To directly verify the effectiveness of the proposed method, we present training dynamics of DRPO with different regularization weights $\lambda$ in Figure 2, where $\lambda$$\infty$ corresponds to the reference method DisCO, which does not employ length reward during training. In terms of performance (first and third figures in Figure 2), we can see that DRPO with smaller $\lambda$ values exhibits marginally worse or comparable performance compared with DisCO, while the second and fourth figures show that smaller $\lambda$ values lead to substantial reductions in response length, with $\lambda=0.1$ reducing length by over 50% relative to $\lambda$$\infty$ (DisCO). These observations demonstrate the effectiveness of DRPO to achieve more concise reasoning while maintaining nearly unchanged training performance. In the following section, we will evaluate the generalization of DRPO to test datasets, compared with other strong efficient reasoning baselines.
Figure 3: Comparison of performance-efficiency trade-off. Left is for fine-tuning 1.5B model and right is for fine-tuning 7B model. Grey lines represent the base model performance before finetuning, with generation length of 4698 for 1.5B model and 4119 for 7B model. Squares denote models trained with reference methods without length penalties (i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO). Triangles denote the models trained by other works.
5.2 Comparison with baselines
In this part, we evaluate the effectiveness of the proposed method on test datasets, compared with existing efficient reasoning baselines.
Trade-off between performance and efficiency. We present the trade-off between performance and efficiency for various methods in Figure 3, where the averaged pass@1 over four math datasets of different difficulty is reported. We observe that our proposed DRPO consistently achieves significantly higher performance than all baselines on finetuning 1.5B and 7B models, including the models trained by other work. Notably, relative to the reference learning method (square marker), the proposed DRPO on finetuning 7B model in Figure 3 (Right) efficiently reduces reasoning length from 3053 to 1502 (51% length reduction) with 2.6% loss of performance via varying $\lambda$ to control the tradeoff, demonstrating the effectiveness of DRPO to reduce reasoning length while preserving the reasoning capability. In contrast, all the efficient reasoning baselines suffer from severe performance degradation when the reasoning length is reduced. For example, RLOO-LP reduces reasoning length from 2975 to 1841 (38% length reduction) but incurs a 7.1% loss in performance on finetuning 7B model, with ALP showing a similar trend. Compared with DRPO, these methods trade off more performance for less reduction in length, highlighting the superiority of our method.
Evaluating Trade-off via AES. Additionally, we directly quantify the effectiveness of our proposed method in enhancing reasoning efficiency with minimal performance drop, using the Accuracy-Efficiency Score (AES). AES is positive when the model reduces output length while maintaining or enhancing accuracy, and negative when accuracy deteriorates. For fair comparison, reference model in AES is each method’s counterpart without length reward, i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO. We present the best AES score for each method in Table 1 and defer detailed results in Appendix A.4. From Table 1, we observe that all the baseline methods exhibit a negative AES score for both 1.5B and 7B models, indicating the inefficiency of existing methods in reducing reasoning length while preserving performance. In contrast, DRPO consistently achieves a positive AES score for finetuning both 1.5B and 7B models, highlighting its capability of improving reasoning efficiency while maintaining performance.
Table 1: Accuracy Efficiency Score (AES) Comparison with Baselines. The best AES score for each method is presented.
Method Pass@1 Length AES
1.5B Model RLOO-LP 0.567 2531 -0.129
ALP 0.606 3494 -0.387
HAPO 0.534 1791 -0.519
DRPO 0.624 1527 0.178
7B Model RLOO-LP 0.692 2649 -0.033
ALP 0.679 2170 -0.134
HAPO 0.596 1717 -1.080
DRPO 0.714 1502 0.249
(a) Performance-efficiency tradeoff on individual datasets with increasing difficulty levels from left to right. (a) is for finetuning 1.5B model and (b) is for finetuning 7B model. Squares denote models trained with reference methods without length penalties (i.e., $\lambda$$\infty$ for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ for ALP, $w=0$ for HAPO).
5.3 Length Reduction for Different Problem Difficulty
In this part, we study the impact of length reduction on questions with varying difficulties. In Figure 5(a), we present the performance of different methods across four math datasets of increasing difficulty, from GSM8K (easy) to AIME (very hard). As shown in Figure 5(a), we observe that: (1) on relatively easy questions like GSM8K, DRPO significantly reduces generation length from 1563 to 356 (77.2% length reduction) for 1.5B model and from 969 to 261 (73.1% length reduction) for 7B model with negligible performance drop (-1.1% for 1.5B model and -0.6% for 7B model). In contrast, all the baselines still bring dramatic performance degradation, indicated by a steep slope. (2) With question difficulty increasing from left to right in Figure 5(a), all the models exhibit longer reasoning length and introduce increasing performance drop, which is reasonable since difficult problems inherently require longer reasoning paths to solve. (3) When comparing Figure 5(a)(a) and (b), we can see that 7B models generally demand shorter reasoning length than 1.5B models and are more resilient to length reduction introduced by DRPO, as evidenced by flatter slopes than 1.5B models. (4) Across various difficulty levels, DRPO consistently achieves a better trade-off between performance and length than all the baselines, demonstrating its effectiveness in guiding efficient reasoning.
We conduct further case studies in Appendix A.5 to analyze the efficient reasoning behavior after DRPO training, compared with DisCO. It is shown that models trained by DRPO preserve existing reflection capability while effectively reducing redundant back-and-forth reasoning.
6 Conclusion and Discussions
In this work, we revealed a key limitation of existing RL-based efficient reasoning methods, that their reliance on group-relative advantages can mislead learning when length rewards are included. To address this, we proposed DRPO, which decouples learning signals of correct and incorrect answers so that length penalties reduce the positive signals for correct reasoning but never reverse them. Our method integrates an optimized positive data distribution under a KL regularization into a discriminative objective, for which we derived a closed-form solution, enabling efficient computation using only on-policy data with importance weighting. Experiments on math reasoning tasks demonstrate significant superiority of our approach, compared with existing efficient reasoning methods.
A limitation of this study is that experiments were confined to 1.5B and 7B models due to computational constraints. Extending the proposed approach to larger models or broader reasoning tasks remains an interesting direction. Besides, while DRPO is designed for efficient reasoning, its formulation is general and can be extended to incorporate other rewards on positive data beyond length, such as process rewards. Another interesting direction is to incorporate mechanisms to adapt $\lambda$ to the difficulty level of questions, larger $\lambda$ for more difficult questions and smaller $\lambda$ for easier questions as implied by our results in Figure 5(a), which we leave for future investigation.
Appendix A Appendix
A.1 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
In this part, we derive the optimal policy $P^{*}_{q}$ that maximizes the length reward with a KL regularization:
$\max_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}r_{l}(o)-\lambda\mathbb{D}_{\text{KL}}(P,\pi_{\text{old}}^{+}(\cdot|q)),$ (8)
where $\lambda>0$ is a regularization parameter, $r_{l}(o)$ denotes length reward of response $o$, $\mathcal{P}$ denotes the set of all probability measures $P$ on correct data given $q$, which are absolutely continuous with respect to $\pi_{\text{old}}^{+}(\cdot|q)$, i.e., $\pi_{\text{old}}^{+}(o|q)=0$ indicates $P(o)=0$.
Following prior work Rafailov et al. [2023], Go et al. [2023], we have
$\max_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}r_{l}(o)-\lambda\mathbb{D}_{\text{KL}}(P,\pi_{\text{old}}^{+}(\cdot|q))$
$=$ $\max_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}r_{l}(o)-\lambda\mathbb{E}_{o\sim P}\log\frac{P(o|q)}{\pi_{\text{old}}^{+}(o|q)}$
$=$ $\max_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}\Big(r_{l}(o)-\lambda\log\frac{P(o|q)}{\pi_{\text{old}}^{+}(o|q)}\Big)$
$=$ $\min_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}\Big(\log\frac{P(o|q)}{\pi_{\text{old}}^{+}(o|q)}-r_{l}(o)/\lambda\Big)$
$=$ $\min_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}\Big(\log\frac{P(o|q)}{\frac{1}{Z(q)}\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)}-\log Z(q)\Big)$ (9)
where $Z(q)=\sum_{o}\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)$ is a partition function, which doesn’t depend on $P$.
Let’s first define $\bar{P}(o|q)=\frac{1}{Z(q)}\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)$. Since $\bar{P}(o|q)\geq 0$ for all $o$ and $\sum_{o}\bar{P}(o|q)=1$, $\bar{P}(o|q)$ is a a valid probability distribution. Thus, we can reformulate (A.1) as:
$\min_{P\in\mathcal{P}}\mathbb{E}_{o\sim P}\Big(\log\frac{P(o|q)}{\frac{1}{Z(q)}\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)}-\log Z(q)\Big)$
$=$ $\min_{P\in\mathcal{P}}\Big(\mathbb{D}_{\text{KL}}(P,\bar{P})-\log Z(q)\Big)$ (10)
Since $Z(q)$ doesn’t depend on $P$, the minimum of (A.1) is achieved by minimizing the first KL term. With Gibbs’ inequality that KL-divergence is minimized at 0 if and only if the two distributions are identical. Therefore, we have the optimal solution:
$P_{q}^{*}(o)=\bar{P}(o|q)$ $=\frac{1}{Z(q)}\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)$ (11)
$=\frac{\pi_{\text{old}}^{+}(o|q)\exp(r_{l}(o)/\lambda)}{\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\exp(r_{l}(o)/\lambda)}.$ (12)
A.2 Detailed Hyperparameter Setting
In this section, we provide detailed hyperparameter settings used in our experiments.
For all the methods, we employ the AdamW optimizer with a weight decay of 0.01 and set constant learning rate to $2\mathrm{e}^{-6}$ for 1.5B model and $1\mathrm{e}^{-6}$ for 7B model, following  Li et al. [2025a]. We set the batch size to 128 for each step of RL, the mini-batch size to 32 for each iteration of model update, and sample 8 responses per question for training. The generation budget is limited to 8k tokens for both training and evaluation. The temperature is set to 0.6 for training.
For RLOO-LP, we use RLOO advantage estimator and clip ratio $\epsilon=0.2$. We tune their weight parameter $\alpha\in\{0.05,0.1,0.2\}$. Following their paper, we normalize its loss by the length of the response.
For ALP, we follow their paper to use GRPO advantage estimator and normalize its loss by the total number of tokens. The KL Coefficient is set to 0.001. We tune their penalty weight $\beta\in\{1\mathrm{e}^{-9},1\mathrm{e}^{-8},1\mathrm{e}^{-7}\}$ and and clip ratio $\epsilon=0.2$
For HAPO, we follow their paper to use GRPO advantage estimator, set the cutoff c=0.7c=-0.7, KL Coefficient to 0, and clip ratio $\epsilon=0.2$. The loss is normalized by the length of the response. We tune their weight parameter $w\in\{0.01,0.1,1\}$.
For the proposed DRPO, we follow  Li et al. [2025a] to set constraint value $\delta=1e^{-4}$, penalty constant $\beta_{0}=1e^{3}$, $\tau=10$. we tune regularization parameter $\lambda\in\{0.5,0.2,0.1\}$.
Table 2: Various reward designs for efficient reasoning, where $r_{c}(o|q)=\mathbb{I}(o\text{ is correct})\in\{1,0\}$ denotes correctness reward.
Method $r(o,q)$
RLOO-LP [Arora and Zanette, ] $r_{c}(o,q)-\alpha*r_{c}(o,q)*\sigma(\frac{|o|-\text{mean}\{|o_{i}|,r_{c}(o_{i},q)=1\}}{\text{std}\{|o_{i}|,r_{c}(o_{i},q)=1\}})$
ALP [Xiang et al., ] $r_{c}(o,q)-\beta*|o|*\max(\text{mean}\{r_{c}(o_{i},q)\},K^{-1})$
HAPO [Huang et al., ] $r_{c}(o,q)+w*\max\Big(\cos(\min(\frac{\pi}{2}\frac{|o|}{h(q)},\pi)),c\Big)r_{c}(o,q)+w*\min\Big(\cos(\min(\frac{\pi}{2}\frac{|o|}{h(q)},\pi)),0\Big)(1-r_{c}(o,q))$
L1-MAX [Aggarwal and Welleck, ] $r_{c}(o,q)*\text{clip}(\alpha(L_{T}-|o|)+\delta),0,1)$
SB [Yi et al., 2025] $\alpha*r_{c}(o,q)-\beta*\text{abs}\Big(|o|-L_{SOL}(q)\Big)$,
where $L_{SOL}(q)=\text{min}\{|o_{i}|,r_{c}(o_{i},q)=1\}$ if at least one response is correct, $\text{mean}\{|o_{i}|\}$ otherwise
LASER-D [Liu et al., 2025b] $r_{c}(o,q)+r_{c}(o,q)*\alpha\mathbb{I}(|o|\leq L_{A})$
Table 3: Illustration of negative learning signal of correct outputs of existing reward designs.
Method Hyperparamter Length Correctness Reward Advantage
RLOO-LP $\alpha$=0.4 [1500, 1200, 1900, 2200, 2800, 2000, 3600, 6400, 1300, 1200] [1, 1, 1, 1, 1,
1, 1, 1, 0, 0]
[0.87, 0.89, 0.85, 0.83, 0.79,
0.84, 0.74, 0.63, 0., 0. ]
[ 0.25, 0.27, 0.23, 0.21, 0.16,
0.22, 0.11, -0.02, -0.72, -0.72]
ALP $\beta$=0.0001 [1500, 1200, 1900, 2200, 2800, 2000, 3600, 6400, 1300, 1200] [1, 1, 1, 1, 1,
1, 1, 1, 0, 0]
[ 0.88, 0.9, 0.85, 0.82, 0.78,
0.84, 0.71, 0.49, -0.1 , -0.1 ]
[ 0.74, 0.8, 0.65, 0.58, 0.46,
0.63, 0.28, -0.32, -1.92, -1.9 ]
HAPO ww=1, cc=-0.7, hh=1200 [1500, 1200, 1900, 2200, 2800, 2000, 3600, 6400, 1300, 1200] [1, 1, 1, 1, 1,
1, 1, 1, 0, 0]
[ 0.62, 1., 0.3, 0.3, 0.3,
0.3, 0.3, 0.3, -0.13, -0. ]
[ 0.99, 2.29, -0.1, -0.1, -0.1,
-0.1, -0.1, -0.1
, -1.57 ,-1.12]
L1-MAX $\alpha$=0.0003, $L_{T}$=4000, $\delta$=0.5 [1500, 1200, 1900, 2200, 2800, 2000, 3600, 6400, 1300, 1200] [1, 1, 1, 1, 1,
1, 1, 1, 0, 0]
[1., 1., 1., 1. , 0.86 ,
1., 0.62 ,0., 0., 0. ]
[0.8, 0.8, 0.8, 0.8 , 0.48,
0.8, -0.06, -1.48, -1.48, -1.48 ]
SB $\alpha$=2, $\beta$=0.001 [1500, 1200, 1900, 2200, 2800, 2000, 3600, 6400, 1300, 1200] [1, 1, 1, 1, 1,
1, 1, 1, 0, 0]
[ 1.7, 2., 1.3, 1., 0.4, ,
1.2, -0.4, -3.2, -0.1, 0.]
[ 0.92, 1.14, 0.64, 0.43, 0.01,
0.57, -0.56, -2.53, -0.35, -0.28]
LASER-D $\alpha$=0.5, $L_{A}$=4000 [1500, 1200, 1900, 2200, 2800, 2000, 3600, 6400, 1300, 1200] [1, 1, 1, 1, 1,
1, 1, 1, 0, 0]
[1.5, 1.5, 1.5, 1.5, 1.5,
1.5, 1.5, 1., 0., 0. ]
[ 0.59, 0.59, 0.59, 0.59, 0.59,
0.59, 0.59, -0.25, -1.94, -1.94]
A.3 Limitation of incorporating length reward with group advantage
In this part, we summarize different reward designs of existing baselines, which incorporate a length reward to encourage efficient reasoning, in Table 2. To illustrate the inherent limitation of incorporating length reward with group advantage, in Table 3, we provide detailed examples of how these reward designs fail to work with group advantage, resulting in misleading learning signals. Specifically, we follow the hyperparameters used in their paper to calculate the advantage with RLOO advantage estimator for RLOO-LP method and GRPO advantage estimator (i.e., Eqn. (2)) for other methods. RLOO advantage estimator is calculated as $A(o_{i}|q)=r(o_{i}|q)-\text{mean}({r(o_{1}|q),\cdots,r(o_{i-1}|q),r(o_{i+1}|q),\cdots,r(o_{G}|q)})$. As indicated by red values in Table 3, all reward designs produce varying amounts of misleading learning signals. We see that HAPO suffers most, yielding incorrect learning directions in 6 out of 10 cases. This could help explain why HAPO exhibits larger performance degradation than other baselines in our experiments.
A.4 Detailed AES performance
In this part, we present detailed AES performance for each method in Table 4 and  5, where bold values denote the best AES performance for each method. We observe that all baseline methods yield negative AES scores for all settings, underscoring their inefficiency in preserving performance while reducing reasoning length. In contrast, DRPO consistently achieves positive AES scores for most cases, demonstrating its effectiveness in improving reasoning efficiency without sacrificing performance.
Table 4: Detailed AES performance for 1.5B models.
Method GSM8K MATH500 OlympiadBench AIME Ave.
RLOO_LP ($\alpha$=0.1) 0.078 0.021 -0.423 -0.738 -0.172
RLOO_LP ($\alpha$=0.2) 0.246 0.325 -0.279 -0.694 -0.129
RLOO_LP ($\alpha$=0.4) 0.026 -0.13 -0.454 -1.18 -0.412
ALP ($\beta$=1e-9) 0.06 0.016 -0.603 -1.571 -0.387
ALP ($\beta$=1e-8) -0.189 -0.088 -0.91 -2.229 -0.602
ALP ($\beta$=1e-7) -1.019 -0.518 -1.887 -3.774 -1.4
HAPO (ww=0.01) -0.209 -0.348 -0.879 0.202 -0.519
HAPO (ww=0.1) -0.74 -0.811 -1.066 -0.457 -0.969
HAPO (ww=1) -1.05 -1.127 -0.841 -0.182 -1.063
DRPO ($\lambda$=0.5) 0.296 0.21 0.152 -0.143 0.093
DRPO ($\lambda$=0.2) 0.614 0.296 0.184 -0.597 0.164
DRPO ($\lambda$=0.1) 0.662 0.332 0.098 -0.729 0.178
Table 5: Detailed AES performance for 7B models.
Method GSM8K MATH500 OlympiadBench AIME Ave.
RLOO-LP ($\alpha$=0.1) 0.391 0.23 -0.051 -0.532 -0.033
RLOO-LP ($\alpha$=0.2) 0.211 0.284 -0.046 -0.829 -0.122
RLOO-LP ($\alpha$=0.4) 0.045 -0.04 -0.103 -0.709 -0.331
ALP ($\beta$=1e-9) 0.451 0.265 -0.297 -0.984 -0.134
ALP $\beta$=1e-8) 0.161 -0.351 -1.048 -1.798 -0.68
ALP ($\beta$=1e-7) -0.923 -1.211 -1.745 -2.679 -1.573
HAPO (ww=0.01) -0.105 -1.001 -1.318 -1.676 -1.08
HAPO (ww=0.1) -0.999 -1.483 -1.252 -1.197 -1.42
HAPO (ww=1) -1.483 -1.407 -1.344 -1.503 -1.6
DRPO ($\lambda$=0.5) 0.053 0.106 -0.126 -0.155 -0.007
DRPO ($\lambda$=0.2) 0.439 0.303 0.015 -0.33 0.115
DRPO ($\lambda$=0.1) 0.672 0.514 0.254 -0.455 0.249
Algorithm 1 Decoupled Reward Policy Optimization (DRPO)
1:Input: Initial policy model $\pi_{0}$, reward function $r$, question set $\mathcal{D}$, hyperparameter $\delta,\beta,\tau,\lambda$.
2: Policy model $\pi_{\theta}=\pi_{0}$
3:for Step $1,\cdots,T$ do
4:  Sample a batch of questions $\mathcal{B}$ from $\mathcal{D}$
5:  Update the old policy model $\pi_{\text{old}}=\pi_{\theta}$
6:  For each question $q\in\mathcal{B}$, sample $n$ responses $\{o_{i}\}_{i=1}^{n}\sim\pi_{\text{old}}(\cdot|q)$ denoted by $S_{q}$ and partition it into $S^{+}_{q}$and $S^{-}_{q}$ based on correctness rewards $r(o_{i}|q)\in\{0,1\}$
7:  for minibatch $\mathcal{B}_{m}\in\mathcal{B}$ do
8:   Compute KL divergence estimator by $\hat{\mathbb{D}}_{KL}(\theta)=\frac{1}{\sum_{q\in\mathcal{B}_{m}}\sum_{o\in S_{q}}|o|}\sum\limits_{q\in\mathcal{B}_{m}}\sum\limits_{o\in S_{q}}\sum\limits_{t=1}^{|o|}\log\frac{\pi_{\theta_{old}}(o_{t}|q,o_{<t})}{\pi_{\theta}(o_{t}|q,o_{<t})}$
[…]
10:   Compute gradient estimator of a penalty function of the constraint by $G_{2}=2\beta_{0}[\hat{\mathbb{D}}_{KL}(\theta)-\delta]_{+}\nabla\hat{\mathbb{D}}_{KL}(\theta)$
11:   Update $\pi_{\theta}$ with Adam-W using the gradient estimator $G=G_{1}+G_{2}$
12:  end for
13:end for
A.5 Case Study
We analyze the reasoning path of DRPO, compared with DisCO, which corresponds to DRPO with $\lambda=+\infty$. Figure 5 shows the reasoning paths on an easy prompt, where DRPO reaches the correct answer with clear reasoning in only 89 tokens, achieving a 6× reduction compared to the 526 tokens required by DisCO. Although DisCO also produces the correct answer, its reasoning is highly redundant, including four occurrences of reflection words (e.g., wait) and three repeated answers. By contrast, DRPO’s reasoning involves only a single reflection and one correct answer.
We further evaluate the reasoning efficiency on a more challenging prompt in Figure 6 and Figure 7. As shown in Figure 6, DisCO requires 4497 tokens, about 10 times longer than the 455 tokens of DRPO, to solve the question. Moreover, the reasoning of DisCO contains 78 reflection words and 19 answers, six of which are incorrect. In comparison, DRPO produces only two correct answers and five reflection words as seen in Figure 7. It is worth noticing that reasoning from DisCO explicitly says that ’but I’m not 100% sure’ before concluding the final answer. It suggests that excessive back-and-forth in the reasoning, indicated by frequent use of reflection words, may result in less confident answers.
Figure 5: Example reasoning for Prompt 1 from DisCo (DRPO $\lambda=+\infty$) and DRPO ($\lambda=0.1$). Words with green color are answers, and with blue color are reflection words. DRPO reaches the correct answer with clear reasoning in only 89 tokens, achieving a 6× reduction compared to the 526 tokens required by DisCO.
Figure 6: Example reasoning for Prompt 2 from DisCo (DRPO $\lambda=+\infty$) . Words in green/red color are correct and incorrect answers, respectively. Blue words are reflection words. The reasoning of DisCO contains 78 reflection words and 19 answers, six of which are incorrect.
Figure 7: Example reasoning for Prompt 2 from DRPO ($\lambda=0.1$). Words in green/red color are correct and incorrect answers, respectively. Blue words are reflection words. DisCO uses 4497 tokens, about 10 times longer than the 455 tokens of DRPO.